OpenStudy (anonymous):

Write the equation of a line that is perpendicular to the given line and that passes through the given point. 2x + 4y = -6; (2,5)

OpenStudy (anonymous):

https://www.desmos.com/calculator Welcome to OpenStudy.

OpenStudy (anonymous):

Thanks for the welcome :3 but that website doesn't help XC here are some of the answers I can pick from A: y= -2x+1 B: y= 1/2x-8 C: y= 2x+1 D: y= 1/2x+1